OpenStudy (anonymous):

What is the product of the quantity m squared plus 2 times m minus 8 all over 6 times p • 12 times p all over 4 times m plus m squared?

OpenStudy (pfenn1):

Is this your expression?\[\frac{m^2+2m-8}{6p} \times \frac{12p}{4m+m^2}\]

OpenStudy (anonymous):

yes it is

OpenStudy (pfenn1):

Can you factor \[m^2+2m-8=(m+a)(m+b)\]where a+b = 2 and ab = -8?

OpenStudy (anonymous):

i see where your going with this i got it thanks lol

OpenStudy (anonymous):

wasnt sure how to start it

OpenStudy (pfenn1):

You're welcome.